OpenStudy (johnw):

lim x->∞ sin(e^x) x ln(x^3+2)+5 / √x Calculate the limit, if it exists.

OpenStudy (johnweldon1993):

Assuming from the post it looks like: \[\large \lim_{x \rightarrow \infty} \frac{\sin(e^x)x\ln(x^3+2)+5}{\sqrt{x}}\] ?

OpenStudy (johnw):

nope, sin(e^x) is separate and is multiplicated with ln(x^3+2)+5/√x

OpenStudy (johnweldon1993):

That's what I thought, the additional 'x' would have just made this diverge Regardless: we are working with \[\large \lim_{x\rightarrow \infty}\frac{\sin(e^x)\ln(x^3+2)+5}{\sqrt{x}}\] What would happen right away if you plugged in a HUGE number here for 'x'? What would the numerator equal? What about the denominator?

OpenStudy (johnw):

Yea, of course, now I see it :)

OpenStudy (johnweldon1993):

Cool :D yeah that's usually the first thing to do is to check what would happen if you just plugged it in

OpenStudy (johnw):

The denominator is correct

OpenStudy (johnweldon1993):

So wait, are you good? Or still need help?

OpenStudy (johnw):

The problem is still not solved so I´ll be needing some help, yes :)

OpenStudy (eliesaab):

Wait a minute. I will help you when my guests leave

OpenStudy (johnw):

Thanks! :)

OpenStudy (eliesaab):

Let me do some steps that will make the problem easier \[ \lim_{x->\infty} \frac {\ln(x)}{\sqrt x}=0\]

OpenStudy (eliesaab):

That is easy to prove using L'Hospital's rulle

OpenStudy (eliesaab):

Near infinity \[ \ln(x^3+3)\approx \ln(x^3)=3 \ln (x)\]

OpenStudy (eliesaab):

Combining the above two steps, one has \[ \frac{\ln(x^3+3)}{\sqrt x} \approx= \frac{3 \ln x}{\sqrt x}\to 0 \] when \( x\to \infty \)

OpenStudy (eliesaab):

Now \[ \left | \frac{ \sin(e^x) \ln(x^3 +2)}{\sqrt x} \right|\le \left |\frac{ \ln(x^3 +2)}{\sqrt x}\right |->0\] By the steps before

OpenStudy (eliesaab):

Your limit is zero since the remaining term \( \frac {5}{\sqrt x}\to 0 \) when x goes to infinity

OpenStudy (eliesaab):

Putting all the steps your original limit will go to zero

OpenStudy (johnw):

Ok, yet another limit is zero? Well, that´s math.... Thank you so much for the help! :)

OpenStudy (eliesaab):

You are welcome